OpenStudy (anonymous):

Consider the function f(x)=x^2+5x+2 and the point A(4,2). f′(x)=........ Find a formula for the slope of a line passing through point A, and an arbitrary point on the function. Your answer should be a formula in terms of x. slope=......... Using your answers from the two previous questions, find the slope(s) of the line(s) through point A that is (are) tangent to f(x). If there is more than one line you can separate your answers with a comma (ex. 5,2 ). slope(s)=

OpenStudy (anonymous):

Find the derivative of f(x), then use the 4 or "x" from the point A and use it as the input of the derivative. So, f'(4)=.... That will yield your slope. Then use the point intercept form to find the tangent line y-y1=m(x-x1) Or do you need help finding the derivative?

OpenStudy (anonymous):

so f(x)= 2x+5 and f'(4)=13? how do i get the slope out of this?

OpenStudy (anonymous):

The derivative is directly related to slope. It is the slope of the tangent line. So f'(x)=2x+5, f'(x)=13. Now use point slope form y-y1=m(x-x1) y-2=13(x-4)

OpenStudy (anonymous):

The derivative yields the tangent slope. So f'(x)=slope of f(x)

OpenStudy (anonymous):

so slope is 13(x-4)+2=26x-104? what about the last part?

OpenStudy (anonymous):

the derivative of any function gives you the slope of that function ie (dy/dx)=tan theta

OpenStudy (anonymous):

basically slope at any point of the curve is equal to the derivative wrt x at that point

OpenStudy (anonymous):

so dy/dx = 2x +5 and at the point (4,2) it will become 13 then use the line equation formula

OpenStudy (anonymous):

the slope of 13 is INCORRECT

OpenStudy (anonymous):

that is the second blank

OpenStudy (anonymous):

take any point on the curve as (x , x^2 + 5x + 2) then the slope will be equal to \[m = \frac{ x^2 + 5x +2 - 2 }{ x -4 } = \frac{ x^2 + 5x }{ x -4 }\]

OpenStudy (anonymous):

katie is this correct?

OpenStudy (anonymous):

let me check suvesh

OpenStudy (anonymous):

thats also incorrect

OpenStudy (anonymous):

but that should be correct because i have directly used the formula for slope of a line between 2 points

OpenStudy (anonymous):

suvesh did i type it correctly? please check in the attachment.

OpenStudy (anonymous):

no just type the last part means x^2 + 5x / x -4 part only

OpenStudy (anonymous):

YES!! survesh that was correct

OpenStudy (anonymous):

do you how to do the last blank?

OpenStudy (anonymous):

answers for the last one is 10,-2

OpenStudy (anonymous):

you just have to solve \[\frac{ x^2 + 5x }{ x - 4} = 2 x + 5\] you will get a quadratic whose roots are 10, -2

OpenStudy (anonymous):

when i type the answer in, it says i need to evaluate f'(x) at these points

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

ok then this is the x-value you have got derivative of f(x) = 2 x +5 answers will be (2 x 10 + 5) and (2 x -2 +5) = 25 and 1

OpenStudy (anonymous):

\[f \prime(x) = 2 x +5 \]

OpenStudy (anonymous):

yes!! you got it again. thank you very much suvesh

OpenStudy (anonymous):

do you know how to do cost function questions?

OpenStudy (anonymous):

nope sorry katie

OpenStudy (anonymous):

no worries. thankyou anyway. see you around. bii

OpenStudy (anonymous):

bye